The shrinkage in length of a certain brand of blue jeans is normally distributed with a mean of 1.1000000000000001 inch and a standard deviation of 0.2 inch. What percent of this brand of jeans will shrink more than 1.5 inches?
The possible answers to this question were 2.279..., 0.433, 1.3, 1.779...

Answer by ewatrrr(24785)

mean of 1.10 inch and a standard deviation of 0.2 inch
p(x > 1.5) = p(z > (1.5-1.10)/.2) = p(z > .4/.2) = normalcdf(2, 100) 100 a placeholder
| Using Caculator..............................
normalcdf(2, 100)= .0228 0r 2.28%